The price of a barrel of crude oil in 1998 was $25. By 2004, the price had increased to $54 per barrel. Assuming oil prices exhibit exponential growth, what year will the price of oil be $123 per barrel?
Answer by checkley77(12844)

TRIED TRUNCATING THE 4.8333 TO 4.8 & IT MAKES IT 26 YEARS INSTEAD OF 25.
AFRAID ROUNDING WONT WORK. SORRY
----------------------------------------------------------------------------
(54-25)/(2004-1998)=29/6=4.8333 is the increase per year.
thus to get to $123 per barrel:
123/4.8333=25 years or 1998+25=2023 the year the price is projected to be $123 per barrel.